Fractions
Rewrite in descending (decreasing)order
3/4,5/7,5/8,7/12,9/14

Answer by FrankM(1040)

3/4 = .75
5/7 = .7143
5/8 = .625
7/12 = .583
9/14 = .6429

It was in order till the last fraction. So

3/4 = .75
5/7 = .7143
9/14 = .6429
5/8 = .625
7/12 = .583